# Understanding Code Understandability Improvements in Code Reviews

### v1

@@ Abstract
Motivation: Code understandability plays a crucial role in software development, as developers spend between 58% and 70% of their time reading source code. Improving code understandability can lead to enhanced productivity and save maintenance costs. Problem: Experimental studies aim to establish what makes code more or less understandable in a controlled setting, but ignore that what makes code easier to understand in the real world also depends on extraneous elements such as developers’ background and project culture and guidelines. Not accounting for the influence of these factors may lead to results that are sound but have little external validity. Objective: We aim to investigate how developers improve code understandability during software development through code review comments. Its basic assumption is that code reviewers are specialists in code quality within a project. Method and Results: We manually analyzed 2,401 code review comments from Java open-source projects on GitHub and found that over 42% of all comments focus on improving code understandability, demonstrating the significance of this aspect in code reviews. We further explored a subset of 385 comments related to code understandability and identified eight categories of code understandability concerns, such as incomplete or inadequate code documentation, bad identifier, and unnecessary code. Among the suggestions to improve code understandability, 83.9% were accepted and integrated into the codebase. Among these, only two (less than 1%) end up being reverted later. We also identified types of patches that improve code understandability, ranging from simple changes (e.g., removing unused code) to more context-dependent improvements (e.g., replacing method calling chain by existing API). Finally, we evaluated the ability of four well-known linters to flag the identified code understandability issues. These linters cover less than 30% of these issues, although some of them could be easily added as new rules. Implications: Our findings motivate and provide practical insight for the construction of tools to make code more understandable, e.g., understandability improvements are rarely reverted and thus can be used as reliable training data for specialized ML-based tools. This is also supported by our dataset, which can be used to train such models. Finally, our findings can also serve as a basis to develop evidence-based code style guides. Data Availability: Our data is publicly available at https://codeupcrc.github.io.
@@ 1 Introduction
Reading code is the main activity of software developers [1], with studies [2, 3] estimating that developers spend between 58% and 70% of their time understanding source code. That means that easy-to-read code can save time and money in software development. Code understandability is the extent to which a developer can comprehend and make sense of source code. It covers a number of aspects, such as naming, commenting, structuring [4]. Empirical studies, e.g., [5, 6, 7, 8, 9], have compared several different ways of writing code to find which factors impact code understandability, e.g., abbreviated vs. complete word identifier names [10].
@@
Laboratory studies on code understandability have as their main advantage the possibility to account for and control extraneous factors and objectively establishing relationships between independent and response variables. However, they are not able to account for elements such as project culture and guidelines, and developers’ background, which directly impact how developers understand code. Not accounting for the influence of these factors may lead to results that are sound but have little external validity. In this paper, our motivation is to unveil what developers care about regarding code understandability and how they improve it, in a real-world setting. Our key insight is that this knowledge is available through publicly-available code reviews in open-source software, where code reviewers act as project specialists in code quality.
@@
Code review is a process in software development where developers read and comment on code written by fellow developers in the same project [11, 12, 13]. In a typical code review scenario, code changes are submitted through pull-requests [^1] In projects hosted on Github., and then read by developers who suggest improvements. Once the concerns are addressed and the improvements are performed, the pull-request is approved for integration into the codebase. These suggestions might concern several code quality aspects, including understandability. In this paper we examine code understandability improvements grounded on real comments about code quality made during code review. More specifically, we investigate (i) what code understandability issues are reported by developers during code review, and (ii) what changes developers apply to address those concerns.
@@
Our methodology is founded on systematic manual analysis of comments written by developers during code review in OSS repositories. We selectd 363 Java open-source projects from GitHub that have an active code review process. In the context of pull requests on GitHub, developers can write general or inline comments. The latter are associated to specific lines of the source code under review. Because we want to identify where the reviewers’ concerns reside and what source code changes were applied to address these concerns, we exclusively focused on inline code review comments in this study222hereafter, we use the term code review comment for simplification. We selected a representative sample of 2,401 code review comments for our study. We first manually analyzed all those code review comments and identified the ones concerning code understandability. From this process, we identified 1,012 code review comments related to code understandability improvements. Then, we analyzed these comments and the corresponding remediation changes that developers made to improve code understandability.
@@
We found that more than 42% of all analyzed code review comments suggest code understandability improvements in scopes such as executable source code, code comments, and text literals. This finding highlights the significance of improving code understandability during code review, as suggested by Bacchelli and Bird [11] and Sadowski et al. [12]. We also analyzed in more depth a random sample of 385 code review comments suggesting code understandability improvements. We identified and labeled the concerns related to code understandability, hereafter referred to as “understandability smells”, that were mentioned by reviewers in our sample of 385 code review comments. We discover eight categories of code understandability smells (e.g., incomplete or inadequate code documentation, bad identifier, unnecessary code) in 16 places in the source code (e.g., Method, Call, Literal, Class). Out of these code reviews comments, 83.9% (323) of the suggestions made by the reviewers were accepted, indicating that developers generally agree with suggestions for improving code understandability. In addition, we perform a fine-grained analysis of actual improvements (the 323 accepted code reviews comments) and identify different types of patches that improve code understandability (e.g., modify an identifier to express the meaning or type of an element, remove non-referred constants and other code elements, add or remove horizontal spacing). While some of these patches can be automatically detected and fixed by linters (e.g., unused variables), others are more challenging to identify as they relate to natural language components (e.g., incomplete code documentation) or are dependent on the context of the project and code style choices (e.g., using boolean expressions instead of equivalent functional methods). Furthermore, 97.2% of the applied patches for code review comments were integrated into the codebase of the project and they are very rarely reverted (less than 1%). The remaining 2.8% of patches had their code removed or replaced with a different solution. Finally, an examination of four popular linters shows that they can detect less than 30% of the analyzed code understandability issues. Some additional cases could be supported by the definition of new rules, whereas others would require them to also interact with natural language text (e.g., to detect typos). Our study provides practical guidance for the construction of tools to make code more understandable, e.g., some improvements that could be automated are currently not available in any of four popular linters, and understandability improvements are rarely reverted and thus can be used as reliable training data for specialized ML-based tools. This is also supported by our dataset, which can be used to train such models. Finally, our findings can serve as a basis to develop evidence-based code style guides. More broadly, this is a step in helping developers to focus on more critical tasks during code review, such as verifying the correctness of code or identifying security vulnerabilities.
@@
Data Availability. The resulting dataset of inline code review comments, which contains a classification of developers’ intention with focus on code understandability improvements and the resulting patches created to address them, is publicly available at https://codeupcrc.github.io.
@@ 2 Background
This section provides a brief overview of the two concepts that underpin our investigation: code reviews and code understandability.
@@ 2.1 Code Review
Figure 1: Inline code review suggestion: replace an if statement by a ternary expression [14].
Code review is an assessment of source code to identify a variety of quality issues, for example, defects, security vulnerabilities, poor understandability, and bad design, among others. In collaborative development platforms, code review takes place when source code changes are proposed to a project by a developer, the submitter. Then, the proposed changes are reviewed by other developers, the reviewers, who provide comments with, for instance, suggestions or requests for improvements that the submitter should apply to the code. Submitters and reviewers interact until the reviewers consider that the proposed code changes are good enough to be integrated into the codebase (else the proposed changes are rejected).
GitHub, the most popular collaborative software development platform [^3] https://github.blog/2023-01-25-100-million-developers-and-counting/, allows code review in the context of pull requests. Pull requests’ reviewers can provide general code review comments in the body of the pull request timeline or inline code review comments that are specific to source code lines of the changed files. Inline comments are about fine-grained, localized reviewers’ concerns, because they explicitly point to specific source code lines where the concerns lie. Figure 1 shows an example of an inline code review comment. The source code lines highlighted in green represent a change proposed to a GitHub project through a pull request. A reviewer wrote an inline comment associated to line 190, suggesting the replacement of the if block by a ternary expression. Here, the reviewer’s concern is explicitly about readability.
@@ 2.2 Code Understandability
Legibility and readability are both important factors for making code easy to read, but understandability is the ultimate goal. Legibility refers to how easily individual code elements can be distinguished from each other, while readability refers to how easily developers can comprehend the code as a whole [4]. Ultimately, understandability encompasses both legibility and readability, and refers to how easily developers can extract relevant information from the source code, including non-functional parts, such as documentation, to perform software development or maintenance tasks [15]. In this work, we are interested in understandability.
In the literature, many studies have evaluated how to improve code legibility and readability [16, 17, 8, 7, 18, 19]. All these studies are based on an informal notion of understandability improvement. A code understandability improvement is a refactoring that, when applied to a program, makes its source code better at communicating to the developer what it does, according to to one of more quality criteria [4].
A basic example of this type of change is renaming. Renaming a program element, in general, is not a functional correction and does not change the program functionality. However, a change to the name of a program element may make it clearer for other humans, by better communicating purpose or rationale. The main motivation for renaming is to improve code understandability, although there are some exceptions where renaming is not motivated by understandability, such as when it matches a pattern of reflection. Moreover, certain aspects of source code are pertinent to understandability by nature, specifically documentation and formatting. Although disregarded by compilers, these elements are deliberately integrated into the source code to assist developers in comprehending the code.
@@
Figure 2: Inline code review suggestion: replace a ternary expression by an if statement [22].
@@
Style guides (a.k.a. code conventions) also are adopted to improve code understandability in the real world. Smith et al. [23] defined that “code conventions are a body of advice on lexical and syntactic aspects of code, aiming to standardize low-level code design under the assumption that such a systematic approach will make code easier to read, understand, and maintain”. However, the recommendations proposed by different style guides are not only about understandability and can also be divergent. For example, while the Google Java Style Guide [^4] https://google.github.io/styleguide/javaguide.html suggests the use of braces even when the block is empty or contains only a single statement (e.g., if (a) {b+c}), the Linux Kernel Coding Style [^5] https://www.kernel.org/doc/html/v4.10/process/coding-style.html disagrees with it for single statements (e.g., if (a) b+c). This suggests that some coding alternatives that make source code more understandable are, to some extent, in the eye of the beholder, e.g., they depend on the experience and context of the developer. However, there is insufficient empirical evidence of which code alternatives are preferred during software development considering code understandability to support the creation of a style guide. In this study, we investigate those code alternatives to further help developers choose the best practices and tool builders to create automated support based on empirical evidence from software development teams. We do so in the context of code review, where developers are quality gatekeepers by reviewing and discussing changes until they consider the code is good to be merged into their projects.
@@ 3.1 Research Questions
Method: We analyzed a representative sample of code review comments to determine the rate at which reviewers suggest code understandability improvements. In this direction, we manually classified 2,401 code review comments selected as explained in Section 3.2, assessing whether each comment pertains to improving code understandability. Furthermore, we categorized the associated scopes of these comments (e.g., documentation, executable source code, string literals). Each code review comment was classified by two authors of this paper. We calculated the Kappa coefficient [24] for assessing the agreement between their evaluation. We found $k=0.59$, which is considered a moderate agreement strength [25]. For every case where there was a disagreement, three or more authors engaged in discussion to reach a final classification. Besides this classification, we also verify the scope to which each comment refers, i.e., where in the source file the reviewer has identified an opportunity to improve understandability (e.g., source code and Javadoc). Additionally, we randomly selected and classified 385 code review comments related to code understandability. We distinguished between explicit comments, which indicate that the improvement suggestion is motivated by code understandability, and implicit comments, where the intention of code understandability improvement is inferred from the comment text, code snippet, and the code improvement itself. We also provided a description of why we considered each code review comment to be related to code understandability. In the sample, we found a total of 50 explicit comments and 335 implicit comments related to code understandability.
@@
Method: We have selected 385 code review comments randomly from the 1,012 reviews that were related to code understandability improvements and were classified in Section 4.1. This sample size of code review comments is a representative sample with a confidence level of 95% and a margin of error of 5% of the unlimited population, according to Cochran’s formula [32]. We then identified the issues that triggered the suggestions for change (the what) and group them into categories that we called understandability smells. For example, in one code review [key=821], the reviewer asked to remove three code elements because they were constants that were not being used, i.e., unnecessary code (the “what”). Also, we identified the type of code element targeted for improvement (the “where”) using Spoon [33], a well-established open-source library for analyzing Java source code. Three authors examined the sample of 385 code review comments. They described the issue presented in each comment and identified where the issue occurred in the code. The same authors then discussed the issues and grouped similar comments together. Afterwards, the groups were refined, and categories were established during meetings involving at least five authors.
@@
We used the nodes of Spoon’s AST to group constructs related to each issue pointed out by code reviewers into categories. For example, we group the nodes CtIf and CtSwitch into the Conditional category [^6] The CtIf node refers to an ’if’ statement in the AST, while the CtSwitch node refers to a ’switch’ statement. Both are conditional statements in the Java language.. We considered the most specific AST node to which an understandability smell is associated to determine where it happens. For example, in a code review [key=123], a reviewer requested using an alternative method invocation to evaluate an expression in an if statement, and we categorize the most specific node identified by Spoon, CtInvocation, as a Call. However, in other scenarios, the understandability smell consists of a set of nodes, and we considered the node that is the closest parent of all other nodes. For instance, in another code review [key=1878], the reviewer asked to rewrite part of the method’s body, and the AST were composed by several nodes such as an invocation and an if statement. In this case, the closest parent of those nodes is a Method, represented by CtMethod in Spoon. We also categorized understandability smells as Method when the reviewer requested renaming the method, adding Javadoc to it, or moving it. In two code review comments where the reviewers pointed out multiple issues in different parts of the source code, we classified the “what” and “where” exclusively based on the location where the code review comment was marked. For example, in a code review comment in the class declaration where the reviewer asked to “remove the Db [from class name] and name attributes from the annotation” [key=1772], the code review comment is at the line of the class declaration and we considered that this (Class) is the place where the issue was pointed out.
@@
Novelty: Brown and Parnin [34] investigated the acceptability of GitHub suggested changes, a special kind of code review comment where reviewers recommend a specific change to the code line or lines. However, as far as we know, no study evaluated how developers addressed to the reviewer’s suggestions for code understandability improvements.
@@
Method: For this RQ, we employed the same set of code review comments that were analyzed for RQ4, i.e., comments about understandability where the recommendation made by the reviewer was accepted and a patch was applied. We manually searched subsequent versions of the files modified in that patch to determine whether the patch has been reverted. We considered that a reversion happens when (i) a reviewer makes a suggestion to improve code understandability, (ii) that suggestion is accepted and a patch is submitted as a consequence, incorporating said suggestion, and (iii) subsequently, a new patch is submitted, undoing the understandability improvement. In this process, we account for renamed files and cases where portions that were affected by the patch were moved to other parts of the system, e.g., a different file. If the latest version of the file still includes the patch that was introduced as a response to the suggestion for improvement, we considered that it was not reverted. If the file as a whole or the element (method, class, etc.) where the modification was performed is deleted, we also considered that the change was not reverted, since we cannot ascertain whether the deletion is connected to said change.
Novelty: Kalliamvakou et al. [37] explored the integration rate of proposed changes in GitHub pull requests into the codebase, while Shimagaki et al. [38] assessed the frequency of commits that were subsequently reverted in project histories. However, a gap remains in the existing literature, as there is no study that specifically examines the adoption and reversion of code understandability improvements within the context of code reviews and the history of the codebase.
@@
Motivation: Linters are widely used in real-world software development [39, 40, 41]. Furthermore, they are capable of detecting a number of typical problems in source code. It is then worth investigating the extent to which they capture the code understandability issues that code reviewers point out. Ideally, reviewers should not have to worry about problems that an automated tool can capture.
@@
We utilized the set of code review comments that were previously analyzed for RQ4, which are about understandability smell instances. For each understandability smell, we manually inspected the descriptions of the rules implemented in linters for finding one that could detect the smell. In the end, we calculated the coverage of understandability smells by computing the number of instances that could be detected by at least one of the linters.
Novelty: Previous work [29] investigated the extent to which SonarQube can detect code understandability issues in code reviews by running this tool. In contrast, in this work we manually analyzed the applicability of every rule implemented by four different linters. The issues reported by linters depend on how they are configured [47], as linters sometimes have contradicting rules, e.g., SonarQube includes a rule Close curly brace and the next ”else” [..] should be on two different lines but also one that ends with …on the same line. A manual analysis can capture these inconsistencies.

### v2

@@ Abstract
Context: Code understandability plays a crucial role in software development, as developers spend between 58% and 70% of their time reading source code. Improving code understandability can lead to enhanced productivity and save maintenance costs. Problem: Experimental studies aim to establish what makes code more or less understandable in a controlled setting, but ignore that what makes code easier to understand in the real world also depends on extraneous elements such as developers’ background and project culture and guidelines. Not accounting for the influence of these factors may lead to results that are sound but have little external validity. Goal: We aim to investigate how developers improve code understandability during software development through code review comments. Our assumption is that code reviewers are specialists in code quality within a project. Method and Results: We manually analyzed 2,401 code review comments from Java open-source projects on GitHub and found that over 42% of all comments focus on improving code understandability, demonstrating the significance of this quality attribute in code reviews. We further explored a subset of 385 comments related to code understandability and identified eight categories of code understandability concerns, such as incomplete or inadequate code documentation, bad identifier, and unnecessary code. Among the suggestions to improve code understandability, 83.9% were accepted and integrated into the codebase. Among these, only two (less than 1%) ended up being reverted. We also identified types of patches that improve code understandability, ranging from simple changes (e.g., removing unused code) to more context-dependent improvements (e.g., replacing method calling chains by existing API). Finally, we investigated the potential coverage of four well-known linters to flag the identified code understandability issues. These linters cover less than 30% of these issues, although some of them could be easily added as new rules. Implications: Our findings motivate and provide practical insight for the construction of tools to make code more understandable, e.g., understandability improvements are rarely reverted and thus can be used as reliable training data for specialized ML-based tools. This is also supported by our dataset, which can be used to train such models. Finally, our findings can also serve as a basis to develop evidence-based code style guides. Data Availability: Our data is publicly available at https://codeupcrc.github.io.
@@ 1 Introduction
Reading code is the main activity of software developers [1], with studies [2, 3] estimating that developers spend between 58% and 70% of their time understanding source code. That means that easy-to-read code can save time and money in software development. Code understandability is the extent to which a developer can comprehend and make sense of source code. It covers a number of aspects, such as naming, commenting, and structuring [4]. Empirical studies, e.g., [5, 6, 7, 8, 9], have compared several different ways of writing code to find which factors impact code understandability, e.g., abbreviated vs. complete word identifier names [10].
@@
Laboratory studies on code understandability have as their main advantage the possibility to account for and control extraneous factors and objectively establish relationships between independent and response variables. However, they are not able to account for elements such as project culture and guidelines, and developers’ background, which directly impact how developers understand code. Not accounting for the influence of these factors may lead to results that are sound but have little external validity. In this paper, our motivation is to unveil what developers care about regarding code understandability and how they improve it, in a real-world setting. Our key insight is that this knowledge is available through publicly available code reviews in open-source software, where code reviewers act as project specialists in code quality.
@@
Code review is a process in software development where developers read and comment on code written by fellow developers in the same project [11, 12, 13]. In a typical code review scenario, code changes are proposed through pull requests (e.g., in projects hosted on GitHub) and then reviewed by developers who suggest improvements. Once the concerns are addressed and the improvements are performed, the pull request is approved for integration into the codebase. These suggestions might concern several code quality aspects, including understandability. In this paper, we examine code understandability improvements grounded on real comments about code quality made during code review. More specifically, we investigate (i) what code understandability issues are reported by developers during code review, and (ii) what changes developers apply to address those concerns.
@@
Our methodology is founded on systematic manual analysis of comments written by developers during code review in OSS repositories. We selected 363 Java open-source projects from GitHub that have an active code review process. In the context of pull requests on GitHub, developers can write general or inline comments. The latter are associated with specific lines of the source code under review. Because we want to identify where the reviewers’ concerns reside and what source code changes were applied to address these concerns, we exclusively focused on inline code review comments in this study [^1] Hereafter, we use the term code review comment for simplification.. We selected a representative sample of 2,401 code review comments for our study. We first manually analyzed all those code review comments and identified the ones concerning code understandability. From this process, we identified 1,012 code review comments related to code understandability improvements. Then, we analyzed these comments and the corresponding remediation changes that developers made to improve code understandability.
@@
We found that more than 42% of all analyzed code review comments suggest code understandability improvements in scopes such as executable source code, code comments, and text literals. This finding highlights the significance of improving code understandability during code review, as suggested by Bacchelli and Bird [11] and Sadowski et al. [12]. We also analyzed in more depth a random sample of 385 code review comments suggesting code understandability improvements. We identified and labeled the concerns related to code understandability, hereafter referred to as “understandability smells”, that were mentioned by reviewers in our sample of 385 code review comments. We discover eight categories of code understandability smells (e.g., incomplete or inadequate code documentation, bad identifier, unnecessary code) in 16 places in the source code (e.g., Method, Call, Literal, Class). Out of these code review comments, 3.9% (323) of the suggestions made by the reviewers were accepted, indicating that developers generally agree with suggestions for improving code understandability. In addition, we performed a fine-grained analysis of actual improvements (the 323 accepted code review comments) and identified different types of patches that improve code understandability (e.g., modify an identifier to express the meaning or type of an element, remove non-referred constants and other code elements, add or remove horizontal spacing). While some of these patches can be automatically detected and fixed by linters (e.g., unused variables), others are more challenging to identify as they relate to natural language components (e.g., incomplete code documentation) or are dependent on the context of the project and code style choices (e.g., using boolean expressions instead of equivalent functional methods). Furthermore, 97.2% of the applied patches for code review comments were integrated into the codebase of the project and they are very rarely reverted (less than 1%). The remaining 2.8% of patches had their code removed or replaced with a different solution. Finally, an examination of four popular linters shows that they can detect less than 30% of the analyzed code understandability issues. Some additional cases could be supported by the definition of new rules, whereas others would require them to also interact with natural language text (e.g., to detect typos). Our study provides practical guidance for the construction of tools to make code more understandable, e.g., some improvements that could be automated are currently not available in any of four popular linters, and understandability improvements are rarely reverted and thus can be used as reliable training data for specialized ML-based tools. This is also supported by our dataset, which can be used to train such models. Finally, our findings can serve as a basis to develop evidence-based code style guides. More broadly, this is a step in helping developers focus on more critical tasks during code review, such as verifying the correctness of code or identifying security vulnerabilities.
@@
Data Availability. The resulting dataset of inline code review comments, which contains a classification of developers’ intention with a focus on code understandability improvements and the resulting patches created to address them, is publicly available at https://codeupcrc.github.io.
@@ 2 Background
This section provides a brief overview of the two concepts that underpin our investigation: code review and code understandability.
@@ 2.1 Code Review
Figure 1: Inline code review suggestion to replace an if statement by a ternary expression [14].
Code review is an assessment of source code to identify a variety of quality issues, such as defects, security vulnerabilities, poor understandability, and bad design. In collaborative development platforms, code review takes place when source code changes are proposed to a project by a developer, the submitter. Then, the proposed changes are reviewed by other developers, the reviewers, who provide comments with, for instance, suggestions or requests for improvements that the submitter should apply to the code. Submitters and reviewers interact until the reviewers consider that the proposed code changes are good enough to be integrated into the codebase (else the proposed changes are rejected).
GitHub, the most popular collaborative software development platform [^2] https://github.blog/2023-01-25-100-million-developers-and-counting/, allows code review in the context of pull requests. Pull request reviewers can provide general code review comments in the body of the pull request timeline or inline code review comments that are specific to source code lines of the changed files. Inline comments are about fine-grained, localized reviewers’ concerns, because they explicitly point to specific source code lines where the concerns lie. Figure 1 shows an example of an inline code review comment. The source code lines highlighted in green represent a change proposed to a GitHub project through a pull request. A reviewer wrote an inline comment associated with line 190, suggesting the replacement of the if block by a ternary expression. In such a case, the reviewer’s concern is explicitly about readability.
@@ 2.2 Code Understandability
Legibility and readability are both important factors for making code easy to read, but understandability is the ultimate goal. Legibility refers to how easily individual code elements can be distinguished from each other, while readability refers to how easily developers can comprehend the code as a whole [4]. Ultimately, understandability encompasses both legibility and readability and refers to how easily developers can extract relevant information from the source code, including non-functional parts, such as documentation, to perform software development or maintenance tasks [15]. In this work, we are interested in understandability.
In the literature, many studies have evaluated how to improve code legibility and readability [16, 17, 8, 7, 18, 19]. All these studies are based on an informal notion of understandability improvement. A code understandability improvement is a refactoring that, when applied to a program, makes its source code better at communicating to the developer what it does, according to one of more quality criteria [4].
A basic example of this type of change is renaming. Renaming a program element, in general, is not a functional correction and does not change the functionality of the program. However, a change to the name of a program element may make it clearer for other humans, by better communicating purpose or rationale. The main motivation for renaming is to improve code understandability, although there are some exceptions where renaming is not motivated by understandability, such as when it matches a reflection pattern. Moreover, certain aspects of source code are pertinent to understandability by nature, specifically documentation and formatting. Although disregarded by compilers, these elements are deliberately integrated into the source code to assist developers in comprehending the code.
@@
Figure 2: Inline code review suggestion to replace a ternary expression by an if statement [22].
@@
Style guides (a.k.a. code conventions) also are adopted to improve code understandability in the real world. Smith et al. [23] defined that “code conventions are a body of advice on lexical and syntactic aspects of code, aiming to standardize low-level code design under the assumption that such a systematic approach will make code easier to read, understand, and maintain”. However, the recommendations proposed by different style guides are not only about understandability and can also be divergent. For example, while the Google Java Style Guide [^3] https://google.github.io/styleguide/javaguide.html suggests the use of braces even when the block is empty or contains only a single statement (e.g., if (a) {b+c}), the Linux Kernel Coding Style [^4] https://www.kernel.org/doc/html/v4.10/process/coding-style.html disagrees with it for single statements (e.g., if (a) b+c). This suggests that some code alternatives that make source code more understandable are, to some extent, in the eye of the beholder, e.g., they depend on the experience and context of the developer. However, there is insufficient empirical evidence of which code alternatives are preferred during software development considering code understandability to support the creation of a style guide. In this study, we investigate those code alternatives to further help developers choose the best practices and tool builders to create automated support based on empirical evidence from software development teams. We do so in the context of code review, where developers are quality gatekeepers who review and discuss changes until they consider the code is good to be merged into their projects.
@@ 3.1 Research Questions
Method: We analyzed a representative sample of code review comments to determine the rate at which reviewers suggest code understandability improvements. In this direction, we manually classified 2,401 code review comments selected as explained in Section 3.2, assessing whether each comment pertains to improving code understandability. Furthermore, we categorized the associated scopes of these comments (e.g., documentation, executable source code, string literals). Each code review comment was classified by two authors of this paper. We calculated the Kappa coefficient [24] to assess the agreement between their evaluation. We found $k=0.59$, which is considered a moderate agreement strength [25]. For every case where there was a disagreement, three or more authors engaged in discussion to reach a final classification. Besides this classification, we also verify the scope to which each comment refers, i.e., where in the source file the reviewer has identified an opportunity to improve understandability (e.g., source code and Javadoc). Additionally, we randomly selected and classified 385 code review comments related to code understandability. We distinguished between explicit comments, which indicate that the improvement suggestion is motivated by code understandability, and implicit comments, where the intention of code understandability improvement is inferred from the comment text, code snippet, and the code improvement itself. We also provided a description of why we considered each code review comment to be related to code understandability. In the sample, we found a total of 50 explicit comments and 335 implicit comments related to code understandability.
@@
Method: We selected 385 code review comments randomly from the 1,012 reviews that were related to code understandability improvements and were classified in Section 4.1. This sample size of code review comments is a representative sample with a confidence level of 95% and a margin of error of 5% of the unlimited population, according to Cochran’s formula [32]. We then identified the issues that triggered the suggestions for change (the what) and grouped them into categories that we called understandability smells. For example, in one code review [key=821], the reviewer asked to remove three code elements because they were constants that were not being used, i.e., unnecessary code (the “what”). Also, we identified the type of code element targeted for improvement (the “where”) using Spoon [33], a well-established open-source library for analyzing Java source code. Three authors examined the sample of 385 code review comments. They described the issue presented in each comment and identified where the issue occurred in the code. The same authors then discussed the issues and grouped similar comments together. Afterwards, the groups were refined, and categories were established during meetings involving at least five authors.
@@
We used the nodes of Spoon’s AST to group constructs related to each issue pointed out by code reviewers into categories. For example, we group the nodes CtIf and CtSwitch into the Conditional category [^5] The CtIf node refers to an ’if’ statement in the AST, while the CtSwitch node refers to a ’switch’ statement. Both are conditional statements in the Java language.. We considered the most specific AST node to which an understandability smell is associated to determine where it happens. For example, in a code review [key=123], a reviewer requested using an alternative method invocation to evaluate an expression in an if statement, and we categorize the most specific node identified by Spoon, CtInvocation, as a Call. However, in other scenarios, the understandability smell consists of a set of nodes, and we considered the node that is the closest parent of all other nodes. For instance, in another code review [key=1878], the reviewer asked to rewrite part of the method’s body, and the AST was composed by several nodes such as an invocation and an if statement. In this case, the closest parent of those nodes is a Method, represented by CtMethod in Spoon. We also categorized understandability smells as Method when the reviewer requested renaming the method, adding Javadoc to it, or moving it. In two code review comments where the reviewers pointed out multiple issues in different parts of the source code, we classified the “what” and “where” exclusively based on the location where the code review comment was marked. For example, in a code review comment in the class declaration where the reviewer asked to “remove the Db [from class name] and name attributes from the annotation” [key=1772], the code review comment is at the line of the class declaration and we considered that this (Class) is the place where the issue was pointed out.
@@
Novelty: Brown and Parnin [34] investigated the acceptability of GitHub suggested changes, a special kind of code review comment where reviewers recommend a specific change to the code line or lines. However, as far as we know, no study evaluated how developers addressed reviewer’s suggestions for code understandability improvements.
@@
Method: For this RQ, we employed the same set of code review comments that were analyzed for RQ4, i.e., comments about understandability where the recommendation made by the reviewer was accepted and a patch was applied. We manually searched subsequent versions of the files modified in that patch to determine whether the patch had been reverted. We considered that a reversion happens when (i) a reviewer suggests improving code understandability, (ii) that suggestion is accepted, and a patch is submitted as a consequence, incorporating said suggestion, and (iii) subsequently, a new patch is submitted, undoing the understandability improvement. In this process, we account for renamed files and cases where portions that were affected by the patch were moved to other parts of the system, e.g., a different file. If the latest version of the file still includes the patch that was introduced as a response to the suggestion for improvement, we considered that it was not reverted. If the file as a whole or the element (method, class, etc.) where the modification was performed is deleted, we also considered that the change was not reverted since we cannot ascertain whether the deletion is connected to said change.
Novelty: Kalliamvakou et al. [37] explored the integration rate of proposed changes in GitHub pull requests into the codebase, while Shimagaki et al. [38] assessed the frequency of commits that were subsequently reverted in project histories. However, a gap remains in the existing literature, as no study specifically examines the adoption and reversion of code understandability improvements within the context of code reviews and the history of the codebase.
@@
Motivation: Linters are widely used in real-world software development [39, 40, 41]. Furthermore, they are capable of detecting several typical problems in source code. It is then worth investigating the extent to which they capture the code understandability issues that code reviewers point out. Ideally, reviewers should not have to worry about problems that an automated tool can capture.
@@
We utilized the set of code review comments that were previously analyzed for RQ4, which are about understandability smell instances. For each understandability smell, we manually inspected the descriptions of the rules implemented in linters to find one that could detect the smell. In the end, we calculated the coverage of understandability smells by computing the number of instances that could be detected by at least one of the linters.
Novelty: Previous work [29] investigated the extent to which SonarQube can detect code understandability issues in code reviews by running this tool. In contrast, we manually analyzed the applicability of every rule implemented by four different linters. The issues reported by linters depend on how they are configured [47], as linters sometimes have contradicting rules, e.g., SonarQube includes a rule Close curly brace and the next ”else” [..] should be on two different lines but also one that ends with …on the same line. A manual analysis can capture these inconsistencies.